OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -1 - 13t. Find the instantaneous velocity at t = 8 by finding the derivative.

OpenStudy (trojanpoem):

You're supposed to differentiate: \[\frac{ d }{ dx } x^{n} = n x^{n-1} \] In your case \[s(t) = -1 - 13t\] \[v = \frac{ ds }{ dt } = -13 \]

OpenStudy (anonymous):

Oh okay. So then d is what again? Been a bit since done a problem like this.